An expensive vacuum system can achieve a pressure as low as 1.00 × 10-8 N/m2 at 21.5 °C. How many atoms are there in a cubic centimeter at this pressure and temperature?
Answer #1

PV = nRT

n = PV/RT

n = (10^-8 x 10^-6)/(8.314 x 294.5) = 4.08 x 10^-18

Number of atoms, N = 6.022 x 10^23 x 4.08 x 10^-18 = 2.46 x 10^6

Comment in case any doubt please rate my answer...